# Log Parsing with Self-Generated In-Context Learning and Self-Correction

### v1

@@ 5.1.1. Accuracy
5.1.2. Robustness
Robustness is also an essential indicator for evaluating the effectiveness of log parsers. The strong robustness implies that log parsers can maintain a stable performance across log data with diverse characteristics, thereby demonstrating superior generalizability (Zhu et al., 2019; Jiang et al., 2023; Le and Zhang, 2023b; Xu et al., 2024b). To compare the robustness of AdaParser with all baselines, we illustrate the distribution of each log parser’s metrics across all datasets using box plots.
The results are shown in Figure 6. AdaParser not only achieves the highest accuracy but also exhibits the smallest variance in performance, as evidenced by its narrowest distribution range. This demonstrates that AdaParser exhibits the strongest robustness in parsing a variety of log data. Specifically, the standard deviations of LILAC for GA, FGA, PA, and FTA are 6.8%, 3.6%, 9.7%, and 11.4%, respectively, outperforming the currently most robust log parser, LILAC, which has standard deviations of 8.8%, 6.2%, 15.2%, and 14.0%. The strong robustness of AdaParser is primarily derived from the extensive pre-trained knowledge in LLMs related to diverse log data from different systems. In addition, the SG-ICL-based parser in AdaParser adapts the LLM to the system-specific characteristics of log datasets and is not subject to historical logs, which can often undermine robustness when dealing with diverse log data.
Answer to RQ1: AdaParser outperforms baseline methods in both message-level metrics and template-level metrics. Furthermore, AdaParser exhibits the strongest robustness when parsing diverse log data from different systems.
5.2. RQ2: Can AdaParser adapt to evolving logs?
@@
Table 2 presents the experimental results. We observe that the performance of both AdaParser and LILAC decreases as the proportion of available logs decreases. Nevertheless, AdaParser consistently outperforms LILAC across all ratios of available logs. Remarkably, even in the zero-shot scenario, AdaParser achieves an accuracy that exceeds all existing baselines. Specifically, AdaParser outperforms LILAC by 29.1% and 44.9% for PA and FTA. This highlights the practicality of AdaParser in real-world evolving systems. Furthermore, AdaParser demonstrates a smaller decline in performance compared to LILAC when faced with evolving logs. Specifically, as the ratio of available logs decreases from 20% to 0%, LILAC’s PA and FTA experience a considerable decrease of 28.5% and 31.2%, respectively. In contrast, AdaParser shows a more modest reduction of 14.3% in PA and 10.5% in FTA. These results emphasize the remarkable adaptability of AdaParser to evolving logs.
Answer to RQ2: AdaParser achieves superior performance compared to LLM-based parsers across various ratios of available logs. Even in the absence of historical logs, AdaParser achieves an accuracy surpassing all existing baselines. This highlights the adaptability of AdaParser.
Table 2. Average accuracy among AdaParser with different ratios of available logs.
@@
5.3. RQ3: How does each design contribute to AdaParser?
@@
The results are shown in Table 2, in which the following observations can be made: (1) Removing SG-ICL or self-correction leads to a decrease across all metrics under various ratios of available logs. For instance, removing self-correction leads to a decrease of 8.6% and 14.6% in PA and FTA under zero-shot scenarios, respectively. This decrease is primarily due to the parsing errors in the templates generated by the LLM, indicating that self-correction can mitigate parsing errors and enhance parsing accuracy. (2) The contribution of our designed components becomes increasingly significant as the ratio of available logs decreases. This is because the LLM acquires task-specific knowledge from sufficient available logs through ICL, enabling it to directly generate correct templates. Conversely, when available logs are scarce or entirely absent, the LLM struggles to generate accurate templates. In these cases, our designed components are fully utilized to enhance parsing accuracy. (3) Incorporating SG-ICL into AdaParser not only enhances accuracy but also efficiency. For example, in zero-shot scenarios, AdaParser with SG-ICL processes each dataset in an average of 673 seconds, compared to 839 seconds without SG-ICL. Moreover, our statistical analysis reveals a substantial decrease in the number of self-corrections required for templates generated by the LLM when equipped with SG-ICL. Specifically, the average number of self-correction per dataset for AdaParser with SG-ICL is 40.1, while that for AdaParser without SG-ICL is 83.7. Therefore, this improvement in efficiency can largely be attributed to SG-ICL’s superior adaption to evolving logs, which reduces the number of generated incorrect templates and the consequent need for self-correction. In conclusion, the ablation study demonstrates that both SG-ICL and self-correction are critical to AdaParser’s performance, particularly under constrained log availability.
Answer to RQ3: Both designs of SG-ICL and self-correction contribute to enhancing AdaParser’s performance, especially in scenarios with limited available logs.
Table 3. The performance of AdaParser with different LLMs under zero-shot scenario (%).
@@
5.4. RQ4: How generalizable is AdaParser with different LLMs?
In this RQ, we assess the performance of AdaParser with different LLMs under zero-shot scenarios, i.e., no historical data is available. Our goal is to determine if the performance gains from AdaParser can be generalized to other LLMs. We extend our evaluation beyond the default GPT-3.5 to include four additional representative and popular LLMs, including two closed-source LLMs (i.e., gemini-1.5-flash (Gemini Team, 2024) and claude-3-sonnet (Anthropic, 2024)) and two open-source LLMs (i.e., DeepSeek-V2-chat (DeepSeek-AI, 2024) and qwen1.5-72b-chat (Bai et al., 2023)).
The results are shown in Table 3. We can observe that AdaParser consistently enhances the performance of the employed LLMs across all metrics by a large margin. On average, all LLMs improve by 5.1%, 7.7%, 19.9%, and 23.6% on GA, FGA, PA, and FTA, respectively. Meanwhile, LLMs with relatively weak performance benefit more from AdaParser, suggesting that AdaParser effectively compensates for deficiencies in backbone models. These results demonstrate that AdaParser can be generally applied to different LLMs, maintaining high accuracy.
@@ 5.4. RQ4: How generalizable is AdaParser with different LLMs?
5.5. RQ5: How efficient is AdaParser in parsing large-scale logs?
Besides effectiveness, efficiency is another critical metric in the practical application of log parsers, given the substantial volume of logs (Wang et al., 2022; Le and Zhang, 2023b; Jiang et al., 2023). In this RQ, we assess the efficiency of AdaParser using the public large-scale Loghub-2.0 datasets as described in Sec. 4.2. To measure AdaParser’s efficiency, we record the time it took to complete the parsing process for AdaParser and baselines on all log datasets. We then calculate the average parsing time for each parser, as depicted in Figure 7.
@@ 5.5. RQ5: How efficient is AdaParser in parsing large-scale logs?
Answer to RQ5: AdaParser demonstrates superior efficiency in processing large-scale logs, outperforming semantic-based log parsers by 3.67 to 6.41 times and achieving comparable performance to the most efficient syntax-based log parsers.
Figure 7. Efficiency of AdaParser and baselines on large-scale datasets.
6. Threats to Validity
@@
Privacy Issue. From the perspective of enterprises, logs are sensitive data, as they often encompass a substantial amount of customer and service information. Employing external LLMs to process internal log data may pose risks to privacy and security issues. Actually, AdaParser is a general framework that can incorporate various LLMs as demonstrated in Sec. 5.4. Users can integrate their own LLMs into AdaParser, thereby avoiding privacy issues.
@@
Existing Log parsers can be categorized into syntax-based and semantic-based log parsers. Syntax-based log parsers leverage manually designed features to extract log templates, which have been widely explored in the past. Specifically, syntax-based log parsers can be further subdivided into three categories. (1) Frequency-based parsers: These log parsers (Vaarandi, 2003; Nagappan and Vouk, 2010; Vaarandi and Pihelgas, 2015; Dai et al., 2020) utilize frequent patterns of token position or n-gram information to distinguish templates and variables in log messages. (2) Clustering-based parsers: These log parsers (Fu et al., 2009; Tang et al., 2011; Shima, 2016; Hamooni et al., 2016) compute similarities between log messages to cluster them into different groups and then extract the constant parts of log messages. (3) Heuristic-based parsers: These log parsers (Du and Li, 2016; He et al., 2017; Jiang et al., 2008; Makanju et al., 2009; Yu et al., 2023b) encode expert domain knowledge into general and effective heuristic rules to identify log templates. Semantic-based log parsers are a recently emerging class of parsers and can achieve higher parsing accuracy by mining semantics from log messages. These log parsers (Liu et al., 2022; Huo et al., 2023b; Li et al., 2023; Le and Zhang, 2023b; Yu et al., 2023a, c; Ma et al., 2024; Wu et al., 2024) typically need labeled log data for model training or tuning. For example, Uniparser (Liu et al., 2022) formulates log parsing as a token classification problem, employing Bi-LSTM models for training. VALB (Li et al., 2023) formulates log parsing as a sequence tagging problem and uses Bi-LSTM-CRF models for training. LogPPT (Le and Zhang, 2023b) tunes a pre-trained language model (e.g., RoBERTa) to perform log parsing.
@@
However, recent studies (Khan et al., 2022; Jiang et al., 2023; Petrescu et al., 2023) have shown that existing log parsers struggle with large-scale and complex log data. Given the strong capacities of Large language models (LLMs) in understanding natural language and solving diverse tasks (Zhao et al., 2023), several studies (Le and Zhang, 2023a; JIANG et al., 2023; Liu et al., 2023b; Xu et al., 2024b) have explored the use of LLMs to improve parsing accuracy. Le et al. (Le and Zhang, 2023a) were pioneers in evaluating the performance of LLMs in log parsing, which demonstrates their potential in performing log parsing. Xu et al. (Xu et al., 2024b) proposed DivLog, a method that utilizes the ICL capability of LLMs to enhance log parsing effectiveness. Jiang et al. (JIANG et al., 2023) further combined the LLM-based parser with an adaptive parsing cache, achieving accurate and efficient LLM-based log parsing. Despite these advancements, existing LLM-based parsers rely solely on the responses generated by LLMs as the final parsed results. This can lead to inaccuracies due to the generation of plausible but incorrect templates or overly broad templates, thus reducing parsing accuracy. Moreover, these methods require selecting demonstrations from a fixed candidate set sampled from extensive historical logs. However, in practice, frequent software updates often result in evolving logs, making sufficient historical log data unavailable. In contrast, our proposed method AdaParser, addresses these limitations by leveraging proposed self-generated ICL and self-correction mechanisms, which significantly enhance both the accuracy and adaptability of LLM-based log parsing.
8. conclusion
@@
This parser proposes AdaParser, an effective and adaptive log parsing framework using LLMs with SG-ICL and self-correction. To facilitate accurate log parsing, AdaParser employs the LLM to self-correct potential parsing errors in the templates it generates. Furthermore, AdaParser maintains a dynamic candidate set composed of previously generated templates as demonstrations to adapt evolving log data. Extensive experiments on public large-scale datasets demonstrate that AdaParser outperforms all state-of-the-art methods, even in zero-shot scenarios. Moreover, when integrated with different LLMs, AdaParser consistently enhances the performance of the utilized LLMs by a large margin. We believe AdaParser would benefit both practitioners and researchers in the field of log analysis.

### v2

@@ V-A1 Accuracy
V-A2 Robustness
Robustness is also an essential indicator for evaluating the effectiveness of log parsers. The strong robustness implies that log parsers can maintain a stable performance across log data with diverse characteristics, thereby demonstrating superior generalizability [46, 28, 14, 26]. To compare the robustness of AdaParser with all baselines, we illustrate the distribution of each log parser’s metrics across all datasets using box plots.
The results are shown in Figure 6. AdaParser not only achieves the highest accuracy but also exhibits the smallest variance in performance, as evidenced by its narrowest distribution range. This demonstrates that AdaParser exhibits the strongest robustness in parsing a variety of log data. Specifically, the standard deviations of AdaParser for GA, FGA, PA, and FTA are 6.8%, 3.6%, 9.7%, and 11.4%, respectively, outperforming the currently most robust log parser, LILAC, which has standard deviations of 8.8%, 6.2%, 15.2%, and 14.0%. The strong robustness of AdaParser is primarily derived from the extensive pre-trained knowledge in LLMs related to diverse log data from different systems. In addition, the SG-ICL-based parser in AdaParser adapts the LLM to the system-specific characteristics of log datasets and is not subject to historical logs, which can often undermine robustness when dealing with diverse log data.
Answer to RQ1: AdaParser outperforms baselines in both message-level metrics and template-level metrics. Furthermore, AdaParser exhibits the strongest robustness when parsing diverse logs from different systems.
V-B RQ2: Can AdaParser adapt to evolving logs?
@@
Table II presents the experimental results. We observe that the performance of both AdaParser and LILAC decreases as the proportion of available logs decreases. Nevertheless, AdaParser consistently outperforms LILAC across all ratios of available logs. Remarkably, even in the zero-shot scenario, AdaParser achieves an accuracy that exceeds all existing baselines. Specifically, AdaParser outperforms LILAC by 29.1% and 44.9% for PA and FTA. This highlights the practicality of AdaParser in real-world evolving systems. Furthermore, AdaParser demonstrates a smaller decline in performance compared to LILAC when faced with evolving logs. Specifically, as the ratio of available logs decreases from 20% to 0%, LILAC’s PA and FTA experience a considerable decrease of 28.5% and 31.2%, respectively. In contrast, AdaParser shows a more modest reduction of 14.3% in PA and 10.5% in FTA. These results emphasize the remarkable adaptability of AdaParser to evolving logs.
Answer to RQ2: AdaParser achieves superior performance compared to LLM-based parsers across various ratios of available logs. Even in the absence of historical logs, AdaParser achieves an accuracy surpassing all baselines, underscoring its remarkable adaptability.
TABLE II: Average accuracy among AdaParser with different ratios of available logs.
@@
V-C RQ3: How does each design contribute to AdaParser?
@@
The results are shown in Table II, in which the following observations can be made: (1) Removing SG-ICL or self-correction leads to a decrease across all metrics under various ratios of available logs. For instance, removing self-correction leads to a decrease of 8.6% and 14.6% in PA and FTA under zero-shot scenarios, respectively. This decrease is primarily due to the parsing errors in the templates generated by the LLM, indicating that self-correction can mitigate parsing errors and enhance parsing accuracy. (2) The contribution of our designed components becomes increasingly significant as the ratio of available logs decreases. This is because the LLM acquires task-specific knowledge from sufficient available logs through ICL, enabling it to directly generate correct templates. Conversely, when available logs are scarce or entirely absent, the LLM struggles to generate accurate templates. In these cases, our designed components are fully utilized to enhance parsing accuracy. (3) Incorporating SG-ICL into AdaParser not only enhances accuracy but also efficiency. For example, in zero-shot scenarios, AdaParser with SG-ICL processes each dataset in an average of 673 seconds, compared to 839 seconds without SG-ICL. Moreover, our statistical analysis reveals a substantial decrease in the number of self-corrections required for templates generated by the LLM when equipped with SG-ICL. Specifically, the average number of self-correction per dataset for AdaParser with SG-ICL is 40.1, while that for AdaParser without SG-ICL is 83.7. Therefore, this improvement in efficiency can largely be attributed to SG-ICL’s superior adaption to evolving logs, which reduces the number of generated incorrect templates and the consequent need for self-correction. In conclusion, the ablation study demonstrates that both SG-ICL and self-correction are critical to AdaParser’s performance, especially in scenarios with limited available logs.
Answer to RQ3: Both designs of SG-ICL and self-correction contribute to enhancing AdaParser’s performance, particularly under limited log availability.
TABLE III: The performance of AdaParser with different LLMs under zero-shot scenario (%).
@@
V-D RQ4: How generalizable is AdaParser with different LLMs?
In this RQ, we assess the performance of AdaParser with different LLMs under zero-shot scenarios, i.e., no historical data is available. Our goal is to determine if the performance gains from AdaParser can be generalized to other LLMs. We extend our evaluation beyond the default GPT-3.5 to include four additional representative and popular LLMs, including two closed-source LLMs (i.e., gemini-1.5-flash [49] and claude-3-sonnet [50]) and two open-source LLMs (i.e., DeepSeek-V2-chat [51] and qwen1.5-72b-chat [52]).
The results are shown in Table III. Base refers to the version of AdaParser without both self-correction and SG-ICL. We can observe that AdaParser consistently enhances the performance of the employed LLMs across all metrics by a large margin. On average, all LLMs improve by 5.1%, 7.7%, 19.9%, and 23.6% on GA, FGA, PA, and FTA, respectively. Meanwhile, LLMs with relatively weak performance benefit more from AdaParser, suggesting that AdaParser effectively compensates for deficiencies in backbone models. These results demonstrate that AdaParser can be generally applied to different LLMs, maintaining high accuracy.
@@ V-D RQ4: How generalizable is AdaParser with different LLMs?
V-E RQ5: How efficient is AdaParser?
Besides effectiveness, efficiency is another critical metric in the practical application of log parsers, given the substantial volume of logs [17, 14, 28]. In this RQ, we assess the efficiency of AdaParser using the public large-scale Loghub-2.0 datasets as described in Sec. IV-B. To measure AdaParser’s efficiency, we record the time it took to complete the parsing process for AdaParser and baselines on all log datasets. We then calculate the average parsing time for each parser, as depicted in Figure 7.
@@ V-E RQ5: How efficient is AdaParser?
Answer to RQ5: AdaParser demonstrates superior efficiency in parsing large-scale logs, outperforming semantic-based log parsers by 3.67 to 6.41 times and achieving comparable performance to the fastest syntax-based log parsers.
Figure 7: Efficiency of AdaParser and baselines on large-scale datasets.
VI Discussion
VI-A Applicability and Cost of AdaParser
As demonstrated in Sec. V-D, AdaParser is compatible with various LLMs, including open-source LLMs such as DeepSeek and Qwen. To address privacy concerns, we recommend deploying these open-source LLMs locally using vLLM [53], a library designed for efficient LLM serving and inference. By simply modifying the api_key and base_url in AdaParser’s configuration file, these LLMs can be seamlessly integrated into AdaParser. Therefore, we believe AdaParser is applicable in real-world scenarios.
AdaParser minimizes the cost of querying LLMs by employing a tree-based parser that reduces redundant queries. Additionally, for each query, since the number of tokens in a single log message is generally small (e.g., tens to hundreds of tokens), the cost of querying LLMs remains negligible. Specifically, we calculate the token usage and cost of AdaParser on 14 benchmark datasets. The total input tokens are 1,221,754 and the total output tokens are 93,618. Given the pricing of \$0.5 per 1M input tokens and \$1.5 per 1M output tokens for gpt-3.5-turbo-0125 [54], the total cost for parsing 3.6M log messages is only \$0.75. Note that the latest GPT-4o mini model, which is smarter and cheaper than GPT-3.5-Turbo, is priced at about one-third of GPT-3.5-Turbo. This means the cost of AdaParser can be further reduced.
VI-B Threats to Validity
@@
Privacy Issue. From the perspective of enterprises, logs are sensitive data because they typically contain extensive amounts of customer and operation information. Utilizing external LLMs to process internal logs can introduce significant privacy and security risks. As detailed in Sec. V-D, AdaParser is a general framework that supports the integration of various LLMs. Users can integrate their own LLMs into AdaParser, thereby avoiding privacy issues.
@@
Existing Log parsers can be categorized into syntax-based and semantic-based log parsers. Syntax-based log parsers leverage human-crafted features to extract log templates, which have been widely explored in the past. Specifically, syntax-based log parsers can be further subdivided into three categories. (1) Frequency-based parsers: These log parsers [55, 10, 56, 12] utilize frequent patterns of token position or n-gram information to distinguish templates and variables in log messages. (2) Clustering-based parsers: These log parsers [9, 57, 58, 59] compute similarities between log messages to cluster them into different groups and then extract the constant parts of log messages. (3) Heuristic-based parsers: These log parsers [60, 11, 47, 61, 62, 63] encode domain knowledge into general and effective heuristic rules to identify log templates. Semantic-based log parsers are a recently emerging class of parsers and can achieve higher parsing accuracy by mining semantics from log messages. These log parsers [13, 15, 16, 14, 31, 18, 27, 64] typically need labeled log data for model training or tuning. For example, Uniparser [13] formulates log parsing as a token classification problem, employing Bi-LSTM models for training. VALB [16] formulates log parsing as a sequence tagging problem and uses Bi-LSTM-CRF models for training. LogPTR [64] formulates log parsing as a text summarization problem and uses a pointer mechanism to copy words from log messages. LogPPT [14] tunes a pre-trained language model (e.g., RoBERTa) to perform log parsing.
@@
However, recent studies [45, 28, 65] have shown that existing log parsers struggle with large-scale and complex log data. Given the strong capacities of Large language models (LLMs) in understanding natural language and solving diverse tasks [20], several studies [24, 25, 19, 26] have explored the use of LLMs to improve parsing accuracy. Le et al. [24] were pioneers in evaluating the performance of LLMs in log parsing, which demonstrates their potential in performing log parsing. Xu et al. [26] proposed DivLog, a method that utilizes the ICL capability of LLMs to enhance log parsing effectiveness. Jiang et al. [25] further combined the LLM-based parser with an adaptive parsing cache, achieving accurate and efficient LLM-based log parsing. Despite these advancements, existing LLM-based parsers rely solely on the responses generated by LLMs as the final parsed results. This can lead to inaccuracies due to the generation of plausible but incorrect templates or overly broad templates, thus reducing parsing accuracy. Moreover, these methods require selecting demonstrations from a fixed candidate set sampled from extensive historical logs. However, in practice, frequent software updates often result in evolving logs, making sufficient historical logs unavailable. In contrast, our method AdaParser addresses these limitations by leveraging the proposed SG-ICL and self-correction mechanism, which significantly enhances both the accuracy and adaptability of LLM-based log parsing.
VIII conclusion
@@
This paper proposes AdaParser, an effective and adaptive log parsing framework using LLMs with SG-ICL and self-correction. To facilitate accurate log parsing, AdaParser employs the LLM to self-correct potential parsing errors in its generated templates. Furthermore, AdaParser maintains a dynamic candidate set composed of previously generated templates as demonstrations to adapt evolving log data. Extensive experiments on public large-scale datasets show that AdaParser outperforms state-of-the-art methods, even in zero-shot scenarios. When integrated with different LLMs, AdaParser consistently enhances the performance of the utilized LLMs by a large margin. We believe AdaParser would benefit both practitioners and researchers in the field of log analysis.